### v2

# Inducing Riesz Bases in $L^{2}$ via Composition Operators

Abstract
Let $C_{h}$ be a composition operator mapping $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$ for some open sets $\Omega_{1},\Omega_{2}\subseteq\mathbb{R}^{n}$. We characterize the mappings $h$ that transform Riesz bases of $L^{2}(\Omega_{1})$ into Riesz bases of $L^{2}(\Omega_{2})$. Restricting our analysis to differentiable mappings, we demonstrate that mappings $h$ that preserve Riesz bases have Jacobian determinants that are bounded away from zero and infinity. We discuss implications of these results for approximation theory, highlighting the potential of using bijective neural networks to construct Riesz bases with favorable approximation properties.
Keywords: Composition Operator. Riesz Basis. Orthonormal Basis. Frame.
Mathematics Subject Classification: Primary 47B33, 42C15.
1 Introduction
Consider the measurable spaces $\bigl(\Omega_{1},\mathcal{B}(\Omega_{1})\bigr)$ and $\bigl(\Omega_{2},\mathcal{B}(\Omega_{2})\bigr)$, where $\Omega_{1},\Omega_{2}\subseteq\mathbb{R}^{n}$ are open subsets and $\mathcal{B}(\Omega_{1}),\ \mathcal{B}(\Omega_{2})$ are their respective Borel $\sigma$-algebras. Endow both spaces with the Lebesgue measure $\mu$ and let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable, non-singular mapping, i. e., $\mu(B)=0$ implies $\mu(h^{-1}(B))=0$ for any $B\in\mathcal{B}(\Omega_{1})$. $h$ induces a linear operator
$C_{h}\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{1})\to L(\Omega_{2}),\qquad C_{h}f\mapsto f\circ h,$
where $L(\Omega_{2})$ denotes the space of all measurable functions on $\Omega_{2}$. If $C_{h}$ acts boundedly from $L^{2}(\Omega_{1})$ to $L^{2}(\Omega_{2})$, we say that $C_{h}$ maps $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$ and refer to it as a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$.
The study of composition operators has a long history in functional and operator analysis, particularly on spaces of analytic functions [4], as well as on Hardy and Bergman spaces [14]. Early work by [21] explored their boundedness, invertibility, and compactness in $L^{p}$ spaces, while [23] extended some of these results to operators mapping between different $L^{p}$ spaces. These foundational studies have laid the groundwork for further investigations into the structural and functional properties of composition operators in broader settings.
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. In this article, we investigate the conditions under which the sequence
$(C_{h}\gamma_{n})_{n=0}^{\infty}$ (1.1)
forms a Riesz basis of $L^{2}(\Omega_{2})$ and characterize its dual, see Theorem˜3.1 and Theorem˜3.2, respectively. Given the potential importance of such induced sequences in optimization problems, we restrict our analysis to the case where $h$ is a differentiable mapping. Our results show that all Riesz bases of the form (1.1) are induced precisely by mappings with finite volume distortion; that is, by mappings whose Jacobian determinants are uniformly bounded above and below, see Corollary˜4.1. To illustrate the practical implications of our study, we present a simple numerical example demonstrating how the mapping $h$ can be modelled by an invertible neural network to construct effecitve Riesz bases of the form (1.1).
Our study is motivated by recent computational advances in nuclear-motion theory [18, 24, 26] and condensed matter physics [27, 25], where parameterized and optimized composition operators have been used to construct bases with improved approxinmation properties. While similar problems have been explored in settings such as weighted Hilbert Bergman spaces [12], spaces of holomorphic functions [13], weighted Hardy spaces in the unit disk [6], and Fock spaces [10], a unified perspective for $L^{2}$ spaces remains unexplored. To address this gap, we rely on Singh’s characterization of bounded and invertible composition operators [22], which extends naturally to composition operators between $L^{2}$ spaces on distinct domains.
2 Preliminaries
A natural measure on $\Omega_{1}$ is the push-forward measure $h_{\#}\mu$, defined by
$h_{\#}\mu(B)\mathrel{\mathop{\ordinarycolon}}=\mu(h^{-1}(B)),\qquad\text{for any }B\in\mathcal{B}(\Omega_{1}).$
Since $h$ is assumed to be non-singular, $h_{\#}\mu\ll\mu$, i. e., $h_{\#}\mu$ is absolutely continuous with respect to $\mu$. We denote by $g_{h}$ the Radon-Nikodym derivative of $h_{\#}\mu$ with respect to $\mu$.
Throughout the manuscript, we call the mapping $h$ injective if there exists a measurable transformation $h_{1}\mathrel{\mathop{\ordinarycolon}}\Omega_{1}\to\Omega_{2}$ such that $h_{1}\circ h$ coincides with the identity mapping on $\Omega_{2}$ up to a set of $\mu$-measure zero. $h_{1}$ is called the left inverse of $h$. Similarly, $h$ is called surjective if there exists a measurable transformation $h_{2}\mathrel{\mathop{\ordinarycolon}}\Omega_{1}\to\Omega_{2}$ such that $h\circ h_{2}$ coincides with the identity mapping on $\Omega_{1}$ up to a set of $\mu$-measure zero. $h_{2}$ is called the right inverse of $h$. Finally, $h$ is said to be bijective if there exists a measurable mapping $h^{-1}\mathrel{\mathop{\ordinarycolon}}\Omega_{1}\to\Omega_{2}$ satisfying
$h^{-1}\circ h(x)=x\qquad\text{for a.e. }x\in\Omega_{2},\ \text{and }h\circ h^{-1}(y)=y\qquad\text{for a.e. }y\in\Omega_{1}.$
Necessary and sufficient conditions for a non-singular mapping $h$ to induce a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$ were characterized by Singh and Manhas [21] for the special case $\Omega_{1}=\Omega_{2}$. Later, Takagi and Yakouchi [23] noted that the same result holds for the more general case $\Omega_{1}\neq\Omega_{2}$.
Theorem 2.1.
A non-singular mapping $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ induces a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$ if and only if $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. In this case, the norm of $C_{h}$ is given by
$\left\|C_{h}\right\|=\left\|g_{h}\right\|_{L^{\infty}(\Omega_{1})}^{1/2}.$
Proof.
See [21, Corollary 2.1.2]. ∎
Without loss of generality, we assume that all functions considered in this work are real-valued. The main question of this work is whether the induced sequence $(\phi_{n})_{n=0}^{\infty}\mathrel{\mathop{\ordinarycolon}}=(C_{h}\gamma_{n})_{n=0}^{\infty}$ is complete in $L^{2}(\Omega_{2})$, i.e., whether there exists for any $f$ in $L^{2}(\Omega_{2})$ a sequence of coefficients $(c_{n}(f))_{n=0}^{\infty}$, $c_{n}\in\mathbb{R}$ for any $n\in\mathbb{N}$, such that
$f=\sum_{n=0}^{\infty}c_{n}(f)\phi_{n}.$
We discuss in our work two notions of completeness. First, that of Riesz bases. Here, the coefficients are unique, and are given by the inner product of $f$ with another Riesz basis $(\widetilde{\phi}_{n})_{n=0}^{\infty}$ that is bi-orthogonal to the original Riesz basis, i. e.,
$\langle\widetilde{\phi}_{n},\phi_{m}\rangle_{\Omega_{2}}=\delta_{nm}\qquad\text{for any }n,m\in\mathbb{N},$
where $\langle.,.\rangle_{\Omega_{2}}$ denotes the inner product on $L^{2}(\Omega_{2})$. Second, that of orthonormal bases. Here, the coefficients are unique and given by the inner product of $f$ with the basis functions.
An important result for our analysis is the fact that these complete sequences can be characterized by operators acting on an orthonormal basis [3]. In particular, let $(\gamma_{n})_{n=0}^{\infty}$ be an orthonormal basis of $L^{2}(\Omega_{1})$. Then
  • The Riesz bases of $L^{2}(\Omega_{2})$ are precisely the sequences $(U\gamma_{n})_{n=0}^{\infty}$, where $U\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{1})\to L^{2}(\Omega_{2})$ is a bounded bijective operator.
  • The orthonormal bases of $L^{2}(\Omega_{2})$ are precisely the sequences $(U\gamma_{n})_{n=0}^{\infty}$, where $U\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{1})\to L^{2}(\Omega_{2})$ is a unitary operator.
3 Inducing Bases via Composition Operators
We begin by citing a key result for our analysis. Although this result was originally proven for the specific case $\Omega_{1}=\Omega_{2}$, it is valid for the more general case $\Omega_{1}\neq\Omega_{2}$ using the same proof.
Lemma 3.1.
Let $C_{h}$ be a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. If $C_{h}$ is invertible, its inverse is a composition operator from $L^{2}(\Omega_{2})$ into $L^{2}(\Omega_{1})$.
Proof.
Since $C_{h}$ is invertible, it follows that $g\mathrel{\mathop{\ordinarycolon}}=C_{h}f$ is a characteristic function if and only if $f$ is a characteristic function, see [21, Corollary 2.2.3]. Denote by $A$ the inverse of $C_{h}$. It follows immediately that $Ag$ is a characteristic function if and only if $f$ is a characteristic function. Since $A$ preserves $L^{2}$ characteristic functions and $\Omega_{1}$ and $\Omega_{2}$ are standard Borel spaces, one can deduce that $A$ is a composition operator, see [21, Theorem 2.1.13]. ∎
We now characterize conditions on $h$ for the induced sequence (1.1) to form a Riesz basis.
Theorem 3.1 (Induced Riesz Basis).
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $h$ is injective and there exists $r>0$ such that $g_{h}\geq r$ $\mu$-a.e. on $\Omega_{1}$.
Proof.
Let $U\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{1})\to L^{2}(\Omega_{1})$ be a unitary operator such that $(\gamma_{n})_{n=0}^{\infty}=(U\phi_{n})_{n=0}^{\infty}$ for some orthonormal basis $(\phi_{n})_{n=0}^{\infty}$ of $L^{2}(\Omega_{1})$.
Let $(C_{h}\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{2})$. This implies that $C_{h}U$ is bijective. Since $U$ is unitary, it immediately follows that $C_{h}=C_{h}UU^{-1}$ is bijective, where $U^{-1}$ denotes the inverse of $U$. Denote the inverse of $C_{h}$ by by $C^{-1}_{h}$. From Lemma˜3.1 it follows that $C^{-1}_{h}\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{2})\to L^{2}(\Omega_{1})$ is a composition operator. There exists a mapping $\zeta$ such that $C^{-1}_{h}=C_{\zeta}$. Since $C_{h}C_{\zeta}f=f$ for any $f$ in $L^{2}(\Omega_{2})$ and $C_{\zeta}C_{h}g=g$ for any $g$ in $L^{2}(\Omega_{1})$, it follows that $h$ is bijective with inverse $h^{-1}=\zeta$. For any $B\in\mathcal{B}(\Omega_{2})$ denote by $\chi_{B}$ the characteristic function and observe that
$\mu(B)$ $=\int_{\Omega_{2}}\chi_{B}\ d\mu$
$=\int_{\Omega_{2}}\chi_{B}\circ h^{-1}\circ h\ d\mu$
$=\int_{\Omega_{1}}\chi_{B}\circ h^{-1}\ g_{h}\ d\mu$
$\leq\|g_{h}\|_{L^{\infty}(\Omega_{1})}\ h_{\#}^{-1}\mu(B)$
$\leq\|g_{h}\|_{L^{\infty}(\Omega_{1})}\ \|g_{h^{-1}}\|_{L^{\infty}(\Omega_{2})}\mu(B),$
where $g_{h^{-1}}$ denotes the Radon-Nikodym derivative of the measure $h_{\#}^{-1}\mu$. Since $C_{h^{-1}}$ is a composition operator from $L^{2}(\Omega_{2})$ into $L^{2}(\Omega_{1})$, it follows that $\|g_{h^{-1}}\|_{L^{\infty}(\Omega_{2})}<\infty$. Further, since $C_{h^{-1}}$ is injective, $g_{h^{-1}}>0$ $\mu$-a.e. on $\Omega_{2}$. Set $r=1/\|g_{h^{-1}}\|_{L^{\infty}(\Omega_{2})}$. It follows immediately that $g_{h}\geq r$ $\mu$-a.e. on $\Omega_{1}$.
Now suppose that $h$ is injective and $g_{h}\geq r$ $\mu$-a.e on $\Omega_{1}$ for some $r>0$. Since $h$ is injective, its preimage preserves measurable sets, that is, $h^{-1}(B)=B$ for all $B\in\mathcal{B}(\Omega_{1})$. As shown in [22, Theorem 2.5], this is necessary and sufficient for the associated composition operator $C_{h}$ to have a dense range in $L^{2}(\Omega_{2})$. Together with the lower bound on $g_{h}$, this implies that $C_{h}$ is surjective, see [22, Theorem 2.3]. To see that $C_{h}$ is injective, we observe that for any $B\in\mathcal{B}(\Omega_{1})$ we have
$h_{\#}\mu(B)$ $=\int_{\Omega_{1}}\chi_{B}\ g_{h}\ d\mu$
$\geq r\ \mu(B).$
It follows that $\mu\ll h_{\#}\mu$. This immediately implies that $C_{h}$ is injective. Since $C_{h}$ is also bounded we conclude that $C_{h}U$ is a bounded bijection. Consequently, $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$. ∎
Remark 3.1.
Combining Theorem˜2.1 and Theorem˜3.1 demonstrates that all Riesz bases of the form $(C_{h}\gamma_{n})_{n=0}^{\infty}$ are induced by mappings $h$ that satisfy the following.
  1. (i)
    $h$ is a non-singular injective mapping.
  2. (ii)
    there exist $r,R>0$ such that $r\leq g_{h}\leq R$ $\mu$-a.e. on $\Omega_{1}$.
We observe that these conditions on $h$ ensure its bijectivity.
As we shall see, a Riesz basis of the form (1.1) and its dual are related by a multiplication operator. We now recall the necessary theory to discuss this relationship.
A measurable mapping $k\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{2}$ induces a linear operator $M_{k}$ defined by
$M_{k}\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{2})\to L(\Omega_{2}),\qquad M_{k}f(x)=k(x)f(x)$
for any function $f$ in $L^{2}(\Omega_{2})$ and any $x$ in $\Omega_{2}$. If $M_{k}$ maps $L^{2}(\Omega_{2})$ into $L^{2}(\Omega_{2})$, it is called a multiplication operator from $L^{2}(\Omega_{2})$ into itself. Necessary and sufficient condition for $M_{k}$ to map $L^{2}(\Omega_{2})$ into itself is that $k$ is essentially bounded [23].
Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping that induces a Riesz basis $(C_{h}\gamma_{n})_{n=0}^{\infty}$ of $L^{2}(\Omega_{2})$. In such a case $h$ is invertible and its inverse $h^{-1}$ defines the push-forward measure $h^{-1}_{\#}\mu$. Denote by $g_{h^{-1}}$ the Radon-Nikodym derivative of $h^{-1}_{\#}\mu$ with respect to $\mu$ and note that $g_{h^{-1}}\in L^{\infty}(\Omega_{2})$. It follows that $M_{g_{h^{-1}}}$ is a bounded multiplication operator from $L^{2}(\Omega_{2})$ into itself.
Theorem 3.2.
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$ and let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping such that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$. Then $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ is the bi-orthogonal dual of $(C_{h}\gamma_{n})_{n=0}^{\infty}$.
Proof.
To prove that the sequence $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis it suffices to show that the operator $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection. $C_{h}$ acts boundedly from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Moreover, $(C_{h}\gamma_{n})_{n=0}^{\infty}$ being a Riesz basis implies that $C_{h}$ is a bijection with inverse $C_{h^{-1}}$. This, in turn, implies that $h^{-1}$ is non-singular and $g_{h^{-1}}$ is bounded $\mu$-a.e. on $\Omega_{2}$, see Lemma˜3.1 and Theorem˜2.1. By Theorem˜3.1 it holds that $g_{h^{-1}}$ is bounded away from zero for a.e. $x\in\Omega_{2}$. By combining these results we deduce that $M_{g_{h^{-1}}}$ is a bounded bijection. Therefore, $M_{g_{h^{-1}}}C_{h}$ is a bounded bijection.
To see that $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ are bi-orthogonal observe that for any $n,m\in\mathbb{N}$ we have
$\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\gamma_{m}\circ h)\ g_{h^{-1}}\ d\mu$ $=\int_{\Omega_{2}}(\gamma_{n}\circ h)\ (\gamma_{m}\circ h)\ dh^{-1}_{\#}\mu$
$=\int_{\Omega_{1}}(\gamma_{n}\circ h\circ h^{-1})\ (\gamma_{m}\circ h\circ h^{-1})\ d\mu$
$=\delta_{nm}.$
It remains to show that the pair $(C_{h}\gamma_{n})_{n=0}^{\infty}$ and $(M_{g_{h^{-1}}}C_{h}\gamma_{n})_{n=0}^{\infty}$ are dual. In other words, it remains to show that for any $f$ in $L^{2}(\Omega_{2})$ we have
$f=\sum_{n=0}^{\infty}\langle f,M_{g_{h^{-1}}}C_{h}\gamma_{n}\rangle_{\Omega_{2}}\ C_{h}\gamma_{n}.$
Note that $C_{h^{-1}}f\in L^{2}(\Omega_{1})$ for any $f\in L^{2}(\Omega_{2})$. Since $(\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{1})$ it follows that
$C_{h^{-1}}f$ $=\sum_{n=0}^{\infty}\ \langle C_{h^{-1}}f,\gamma_{n}\rangle_{\Omega_{1}}\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{1}}f\circ h^{-1}\ \gamma_{n}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \biggl(\int_{\Omega_{2}}f\ \gamma_{n}\circ h\ g_{h^{-1}}\ d\mu\biggr)\ \gamma_{n}$
$=\sum_{n=0}^{\infty}\ \langle f,M_{g_{h^{-1}}}C_{h}\gamma_{n}\rangle_{\Omega_{2}}\ \gamma_{n}.$
The result follows by applying $C_{h}$ to both sides. ∎
For the completeness of our results, we conclude this section by remarks on the possibility of generating orthonormal bases and frames of the form (1.1).
Theorem 3.3 (Induced Orthonormal Bases).
Let $(\gamma_{n})_{n=0}^{\infty}$ be an orthonormal basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be a measurable mapping that induces a composition operator $C_{h}$ from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$, i. e., such that $g_{h}$ is bounded $\mu$-a.e. on $\Omega_{1}$. The sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{2})$ if and only if $g_{h}=1$ $\mu$-a.e. on $\Omega_{1}$.
Proof.
Trivial. ∎
Theorem˜3.3 demonstrates that non-trivial orthonormal bases cannot be generated by composition operators alone. Essentially, the problem here is the difficulty of preserving the orthonormality of the underlying basis. To see this, note, e. g., that for any $m\in\mathbb{N}$,
$\int_{\Omega_{2}}(C_{h}\gamma_{m})^{2}\ d\mu=\int_{\Omega_{1}}\gamma_{m}^{2}\ g_{h}\ d\mu.$
Requiring that this equals 1 indeed poses a severe restriction on $h$, namely that $g_{h}=1$ for a.e. $x\in\Omega_{1}$. Generating orthonormal bases of $L^{2}(\Omega_{2})$ can be achieved by using a different approach, such as a weighted composition operator. To this end, assume that $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ is non-singular and bijective. Since it is bijective, it follows that $h(B)\in\mathcal{B}(\Omega_{1})$ for any $B$ in $\mathcal{B}(\Omega_{2})$ [8, Corollary 15.2]. Note that $h$ induces the pullback measure $h^{\#}\mu$ on $\Omega_{2}$ defined by $h^{\#}\mu(B)\mathrel{\mathop{\ordinarycolon}}=\mu(h(B))$ for any $B\in\mathcal{B}(\Omega_{2})$. Denote by $k_{h}$ the Radon-Nikodym derivative of $h^{\#}\mu$ with respect to $\mu$ and note that the operator
$W_{h}f\mathrel{\mathop{\ordinarycolon}}L^{2}(\Omega_{1})\to L(\Omega_{2}),\qquad W_{h}\mathrel{\mathop{\ordinarycolon}}=(f\circ h)\cdot k^{1/2}_{h}$
maps $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Moreover, $W_{h}$ is unitary and hence it follows that $(W_{h}\gamma_{n})_{n=0}^{\infty}$ is an orthonormal basis of $L^{2}(\Omega_{2})$.
We note that sequences of the form $(W_{h}\gamma_{n})_{n=0}^{\infty}$ appeared in some previous studies [19, 17]. However, these results assumed more regularity on $h$.
Remark 3.2 (Flow-Induced Frames).
Another important type of complete sequences in $L^{2}$ is that of frames. In contrast to orthonormal and Riesz bases, expansion coefficients of frames are not necessarily unique. One might wonder what the precise conditions on $h$ are for the induced sequence (1.1) to form a frame. It can be demonstrated that frames of the form $(C_{h}\gamma_{n})_{n=0}^{\infty}$ are induced precisely by non-singular mappings $h$ that are injective and such that $g_{h}$ is bounded away from zero almost everywhere on its support. Deriving a canonical dual frame for such a frame is, however, not as straightforward as in the case of Riesz bases of the form (1.1), since $h$ is not necessarily bijective.
4 Induced Bases via Differentiable Mappings
Bases of the form (1.1) are of interest in optimization problems, particularly those involving differential operators. To facilitate the use of such bases in applications, we state in this section the necessary and sufficient conditions for completeness in $L^{2}(\Omega_{2})$ when $h$ is a differentiable mapping.
Under the conditions of Theorem˜2.1 it immediately follows that $h_{\#}\mu$ is locally finite. Hence, assuming that $h$ is everywhere differentiable, it follows by Lebesgue differentiation theorem [20] that
$g_{h}=\frac{1}{|\det J_{h}|\circ h^{-1}}\hskip 18.49988pt\mu-\text{a.e. on }\Omega_{1},$
where $J_{h}$ denotes the Jacobian of $h$. Using the inverse function theorem, we have
$\frac{1}{|\det J_{h}|\circ h^{-1}}=|\det J_{h^{-1}}|\hskip 18.49988pt\mu-\text{a.e. on }\Omega_{1}.$
Based on the gained regularity we can now state the following.
Corollary 4.1.
Let $(\gamma_{n})_{n=0}^{\infty}$ be a Riesz basis of $L^{2}(\Omega_{1})$. Let $h\mathrel{\mathop{\ordinarycolon}}\Omega_{2}\to\Omega_{1}$ be an everywhere differentiable and injective mapping that induces a composition operator from $L^{2}(\Omega_{1})$ into $L^{2}(\Omega_{2})$. Then $(C_{h}\gamma_{n})_{n={0}}^{\infty}$ is a Riesz basis of $L^{2}(\Omega_{2})$ if and only if $r\leq\det J_{h}\leq R$ for a.e. $x$ in $\Omega_{2}$ and some $r,R>0$.
Proof.
Note that when $h$ is differentiable and injective, the lower and upper bound on $\det J_{h}$ imply that $h$ is bijective. Therefore, the result follows immediately from Theorem˜3.1, and the application of the inverse function theorem. ∎
Observe that in the case where $(C_{h}\gamma_{n})_{n=0}^{\infty}$ is a Riesz basis, its dual is given by
$(M_{\det J_{h}}C_{h}\gamma_{n})_{n=0}^{\infty}=(\gamma_{n}\circ h\ \det J_{h})_{n=0}^{\infty}.$
In one dimension, the condition $r\leq\det J_{h}\leq R$ a.e. on $\Omega_{2}$ simplifies to $r\leq h^{\prime}\leq R$, which means that $h$ is $(r,R)$-bi-Lipschitz, i. e.,
$r\,|x-y|\leq|h(x)-h(y)|\leq R\,|x-y|,\hskip 18.49988pt\text{ for a.e. }\,x,y\in\Omega_{2}.$
Remark 4.1.
Often, continuous differentiability is assumed to use the inverse function theorem. However, differentiability everywhere is enough, see [16, Theorem 1].
5 Application: learning a basis with improved approximation properties
The use of bases for approximation problems is ubiquitous in applied and engineering sciences. Typically, an unknown function $f$ on a domain $\Omega\subseteq\mathbb{R}^{n}$ is approximated in the linear span of a truncated basis $(\gamma_{n})_{n=0}^{N-1}$ of the space $L^{2}(\Omega)$ for some $N\in\mathbb{N}$. Due to the completeness of $(\gamma_{n})_{n=0}^{\infty}$ in $L^{2}(\Omega)$, the approximation error converges to zero as $N\to\infty$. In practice, the accuracy of the approximation and the convergence rate highly depend on the choice of the basis [7]. This dependence renders methods to choose an optimal basis for a given problem highly desirable.
We argue here that one can construct a problem-specific basis by optimizing a perturbation to an orthonormal basis via a composition operator. The following example suggests that a perturbation induced by a function as simple as a shifting mapping can lead to a significant improvement in the approximation of a target function.
Example 5.1.
Let $\Omega_{2}=\Omega_{1}=\mathbb{R}$ and set $(\gamma_{n})_{n=0}^{\infty}$ to be the sequence of Hermite functions. These functions are given by $\gamma_{n}(x)=a_{n}h_{n}(x)\exp(-x^{2}/2)$, where $h_{n}$ is the $n$-th Hermite polynomial and $a_{n}$ is a normalization constant. Let $f$ be a normalized Gaussian function centered around a point $a$, i. e.,
$f(x)=\frac{1}{{\pi}^{1/4}}\exp(-(x-a)^{2}/2),$
for $x\in\mathbb{R}$. Consider approximating $f$ in the linear span of $(\gamma_{n})_{n=0}^{N-1}$ for some $N\in\mathbb{N}$. Hermite functions form an orthonormal basis of $L^{2}(\mathbb{R})$. Therefore, the approximation error in $L^{2}(\mathbb{R})$ converges to zero as $N\to\infty$. However, the approximation error in $L^{2}(\mathbb{R})$ becomes small for $N\gg\frac{e}{2}a^{2}$, where $e$ is Euler’s number.This follows from standard results on Hermite function approximation combined with Stirling’s asymptotic formula for N! [11, Chapter III. Theorem 1.2.]. In other words, to have a good approximation one requires a number of functions that depends nonlinearly on the center of the Gaussian $a$. Consider now the modified basis $(C_{h}\gamma_{n})_{n=0}^{\infty}$ where h(x)=xah(x)=x-a. Note that $f=\gamma_{0}\circ h$, i. e., one needs only one function of the sequence $(C_{h}\gamma_{n})_{n=0}^{\infty}$ to reproduce the target function exactly.
Composing orthonormal bases with linear mappings is a common practice in computational mathematics [7]. Such mappings are often chosen based on insight or formal analytical reasoning [2]. Composition with non-linear mappings are more scarce in the literature, since it is harder to identify effective nonlinear mappings based on insight. However, given a class $\mathfrak{H}$ of adequate mappings, one can use numerical optimization techniques to choose an optimal mapping $h^{*}\in\mathfrak{H}$ for a given problem. In general, one can end up with a mapping $h^{*}$ that leads to a sequence $(C_{h^{*}}\gamma_{n})_{n=0}^{\infty}$ that is not complete. Our results demonstrate that this can be avoided by setting $\mathfrak{H}$ to be a class of invertible mappings with bounded volume distortion.
While there are several ways to construct such classes $\mathfrak{H}$, one popular approach is to use neural networks. Indeed, there exist many invertible neural-network architectures [9], such as, e. g., invertible residual neural networks [1]. Here, the network $h$ can be written as a composition of a sequence of $M$ blocks, i. e., $h=h_{1}\circ h_{2}\circ\ldots\circ h_{M}$, where each block $h_{i}$ is given by
$h_{i}(x)=x+\text{NN}(x),$
and NN is a standard multi-layer perceptron. Requiring NN to be Lipschitz with a Lipschitz constant $L<1$ ensures that $h_{i}$ is a $\bigl((1-L)-(1+L)\bigr)$-bi-Lipschitz. This, in turn, implies that Jacobian determinant of $h$ is bounded from above and below. Such networks are often used to generate complex probability measures from simple ones and are called normalizing flows [15]. Our results suggest that the same flows can be used to generate problem-specific bases. The following numerical example illustrates the idea.
Example 5.2 (Numerical Evidence for Improved Approximation using Composition with a nonlinear Mapping).
Consider the univariate functions $f_{1}$ and $f_{2}$ given by
$f_{1}(x)$ $=\sin(2|x|)\ \exp(-x^{2}/2),$
$f_{2}(x)$ $=\sin(2x)\ \exp(-x^{4}/2)$ (5.1)
and plotted in Figure˜1.
Figure 1: Plot of the functions $f_{1}$ (panel a) and $f_{2}$ (panel b) given by (5.2).
Clearly, the functions $f_{1}$ and $f_{2}$ belong to $L^{2}(\mathbb{R})$ and hence they can be approximated in the linear span of Hermite functions $(\gamma_{n})_{n=0}^{N-1}$ for some $N\in\mathbb{N}$. However, Hermite approximations of both functions converge in a suboptimal manner. For $f_{1}$, this is due to the non-smooth behavior of at $x=0$. For $f_{2}$, this is due to its faster decay with respect to the weight $\exp(-x^{2}/2)$ inherent to Hermite functions.
Consider now approximating $f_{1}$ and $f_{2}$ in the linear span of $(\gamma_{n}\circ h_{1})_{n=0}^{N-1}$ and $(\gamma_{n}\circ h_{2})_{n=0}^{N-1}$, respectively, where $h_{1}$ and $h_{2}$ are bi-Lipschitz mapping that we modeled using a linear mapping composed with an invertible residual neural network of one block, i. e., $h_{1}$ and $h_{2}$ are given by
$h_{1}(x)$ $=\alpha_{1}(x+\text{NN}_{\theta_{1}}(x))+\beta_{1},$
$h_{2}(x)$ $=\alpha_{2}(x+\text{NN}_{\theta_{2}}(x))+\beta_{2}$
where $\text{NN}_{\theta}$ denotes a multi-layer perceptron of one layer composed of 8 hidden units that use nonlinear activation functions. The parameters $\theta_{1},\theta_{2}$ along with the linear parameters $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}\in\mathbb{R}$ were optimized to minimize the $L^{2}$-error in approximating $f_{1}$ and $f_{2}$ in the linear span of $(\gamma_{n}\circ h_{1})_{n=0}^{9}$ and $(\gamma_{n}\circ h_{2})_{n=0}^{9}$, respectively. Figure˜2 shows the perturbing functions $h_{1}$ and $h_{2}$. We estimated the Lipschitz constants of $h_{1}$ and $h_{2}$ by computing the minimum and maximum of $h^{\prime}$ over a uniform grid of $5000$ points discretizing the domain [8,8][-8,8]. The Lipschitz constants are summarized in Table˜1.
Figure 2: Plot of the bi-Lipschitz mappings $h_{1}$ and $h_{2}$ used to perturb Hermite functions for approximating the target functions $f_{1}$ and $f_{2}$.
Table 1: Estimated lower and upper Lipschitz constants for $h_{1}$ and $h_{2}$.
Function $L_{\text{lower}}$ $L_{\text{upper}}$
$h_{1}$ 1.39 2.68
$h_{2}$ 1.61 3.22
Following our theoretical analysis, $(\gamma_{n}\circ h_{1})_{n=0}^{\infty}$ and $(\gamma_{n}\circ h_{2})_{n=0}^{\infty}$ are bases of $L^{2}(\mathbb{R})$ and the duals are given by $\bigl(\gamma_{n}\circ h_{1}\ (h_{1}^{\prime})\bigr)_{n=0}^{N-1}$, $\bigl(\gamma_{n}\circ h_{2}\ (h_{2}^{\prime})\bigr)_{n=0}^{N-1}$, respectively. Using these results, we computed the expansion coefficients for multiple values of $N$ and compared the errors of the resulting approximations with the errors of approximations obtained by using Hermite functions. Figure˜3 shows the approximation error for different values of $N$. The results show that more accurate approximations can be obtained with our choice of the perturbing functions $h_{1}$ and $h_{2}$. Notably, the approximation of $f_{2}$ in the linear span of $(\gamma_{n}\circ h_{2})_{n=0}^{N-1}$ exhibits a larger convergence rate in comparison to the rate achieved by Hermite approximations. No such effect was observed for approximations of $f_{1}$.
Figure 3: Convergence of the $L^{2}$-error in approximating the functions $f_{1}$ (panel a) and $f_{2}$ (panel b) in the linear span of Hermite functions and the perturbed bases.
Figure˜4 shows how Hermite basis functions transform under the composition operators $C_{h_{1}}$ and $C_{h_{2}}$. Clearly, the perturbed basis functions $(\gamma_{n}\circ h_{1})_{n=0}^{3}$ exhibit sharper behavior around the origin, at which the function $f_{1}$ is non-smooth. Similarly, the functions $(\gamma_{n}\circ h_{2})_{n=0}^{3}$ decay faster than Hermite functions to match the fast decay of $f_{2}$.
Figure 4: Plotted are the Hermite functions $(\gamma_{n})_{n=0}^{3}$ as solid black lines, along with their perturbations $(\gamma_{n}\circ h_{1})_{n=0}^{3}$ (red lines with triangle markers) and $(\gamma_{n}\circ h_{2})_{n=0}^{3}$ (blue lines with circle markers). The functions corresponding to n=0, 1, 2, 3n=0,\ 1,\ 2,\ 3 are plotted in panels a, b, c, d, respectively.
In summary, we showed in the previous examples that the approximation properties of Hermite functions can be improved by composing them with a proper linear or nonlinear bi-Lipschitz function. This result is, indeed, not restricted to Hermite functions and holds for other bases. Further, we demonstrated that such bi-Lipschitz mappings can be constructed via normalizing flows, i. e., invertible neural networks. While the examples are rather simple, we note that a few recent works have already made important first steps into more complex applications. Specifically, [5] and [18] utilized bi-Lipschitz normalizing flows to optimize non-linear perturbations to orthonormal bases of $L^{2}$ for solving differential equations. [18] reported several orders-of-magnitude increased accuracy and improved convergence rates upon optimizing a perturbation of the underlying basis. Our results serve as a theoretical framework for such applications and open up new avenues for the design and numerical analysis of problem-specific complete sequences.
Acknowledgments
The authors would like to thank Eleonora Ficola, Stephan Wojtowytsch, Álvaro Fernández Corral and Emil Vogt for useful discussions and constructive feedback.
We acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG) within the Research Training Group GRK 2583 “Modeling, Simulation and Optimization of Fluid Dynamic Applications”.
Author Contributions
Y. S. conceptualized the study, developed the theory, wrote the manuscript and performed the numerical experiments. A.I. contributed to the discussion and interpretation of results, and proofreading of the manuscript.
Data Availability
The manuscript has no associated data.
Code Availability
The codes we developed to produce the numerical results in Section˜5 are available at https://github.com/robochimps/bases_by_composition.
Declarations
Conflict of interest. The authors declare that they have no conflict of interest.